# thanks: Corresponding author

### v1

@@ II.3 GW detection and Fisher analysis
$\Delta{d_{\mathrm{L}}}$ $=\sqrt{\left(\Delta{d_{\mathrm{L}}^{\rm inst}}\right)^{2}+\left(\Delta{d_{\mathrm{L}}^{\rm lens}}\right)^{2}}.$ (23)
It should be added that we assume the distance to each pulsar has been precisely determined. Under this assumption, the derivatives of GW frequencies can be obtained, and the chirp masses of the SMBHBs can be further measured. This enables us to break the degeneracy between ${M_{\rm c}}$ and $d_{\rm L}$, and therefore precisely infer $d_{\rm L}$ from GW amplitudes. The distributions of the $d_{\rm L}$ relative errors and the angular resolutions of simulated SMBHBs versus SNR based on 20, 100, and 200 MSPs with $\sigma_{t}$ = 20 $\rm ns$ and 10-year observation are displayed in Fig. 3. The sky locations of MSPs are randomly selected from the Australia Telescope National Facility pulsar catalog Manchester et al. (2005).
@@
II.4 Inference of the Hubble constant
@@
$H(z)=H_{0}\sqrt{\Omega_{\rm m}(1+z)^{3}+1-\Omega_{\rm m}},$ (24)
@@
$d_{\rm L}(z)=c(1+z)\int_{0}^{z}\frac{1}{H(z^{\prime})}\mathrm{d}z^{\prime},$ (25)
where $c$ is the speed of light in vacuum. Throughout this work, we adopt the flat $\Lambda$CDM model to generate the dark siren data with the best-fit values of $\Omega_{\rm m}=0.3111$ and $H_{0}=67.66\rm~{}km~{}s^{-1}~{}Mpc^{-1}$ given by the Planck 2018 TT, TE, EE+lowE+lensing+BAO data Aghanim et al. (2020). Based on the flat $\Lambda$CDM model, we utilize the GW signals from SMBHBs as dark sirens to infer $H_{0}$, fixing the other cosmological parameters to the Planck 2018 results.
@@
$p\left(H_{0}\mid\mathcal{D}_{\mathrm{GW}},\mathcal{D}_{\mathrm{EM}}\right)$ (26)
$=\frac{\prod_{i=1}^{N_{\mathrm{GW}}}p\left(\mathcal{D}_{\mathrm{GW},i},\mathcal{D}_{\mathrm{EM},i}\mid{H_{0}}\right)p({H_{0}})}{\gamma(H_{0})},$
@@
$p\left(\mathcal{D}_{\mathrm{GW},i},\mathcal{D}_{\mathrm{EM},i}\mid H_{0}\right)$ (27)
@@
$=\int p\left(\mathcal{D}_{\mathrm{GW},i}\mid{d}_{\rm{L}},\alpha,\delta\right)p\left(\mathcal{D}_{\mathrm{EM},i}\mid z,\alpha,\delta\right)p\left({d}_{\rm{L}}\mid z,H_{0}\right)$
×$\qquad\times p_{0}\left(z,\alpha,\delta\mid H_{0}\right)\mathrm{d}{d}_{\rm{L}}\mathrm{d}\alpha\mathrm{d}\delta\mathrm{d}z$
$=\int p\left(\mathcal{D}_{\mathrm{GW},i}\mid{d}_{\rm{L}},\alpha,\delta\right)p\left(\mathcal{D}_{\mathrm{EM},i}\mid z,\alpha,\delta\right)$
×$\qquad\times\delta\left.(d_{\rm{L}}-\hat{d}_{\rm{L}}\left(z,H_{0}\right)\right.)p_{0}\left(z,\alpha,\delta\mid H_{0}\right)\mathrm{d}d_{\rm{L}}\mathrm{d}\alpha\mathrm{d}\delta\mathrm{d}z$
$=\int p\left.(\mathcal{D}_{\mathrm{GW},i}\mid\hat{d}_{\rm L}\left(z,H_{0}\right),\alpha,\delta\right.)p\left(\mathcal{D}_{\mathrm{EM},i}\mid z,\alpha,\delta\right)$
×$\qquad\times p_{0}\left(z,\alpha,\delta\mid H_{0}\right)\mathrm{d}\alpha\mathrm{d}\delta\mathrm{d}z.$
@@
$p\left.(\mathcal{D}_{\mathrm{GW},i}\mid\hat{d}_{\mathrm{L}}\left(z,H_{0}\right),\alpha,\delta\right.)\propto e^{-\chi^{2}/2},$ (28)
with $\chi^{2}=(x-x_{\rm gw})^{T}\mathcal{C}^{-1}(x-x_{\rm gw})$, where $\mathcal{C}$ is the $3\times 3$ covariance matrix only relevant to $(d_{\rm{L}},\alpha,\delta)$, obtained from the Fisher information matrix. $x=(\hat{d}_{\rm{L}}(z,H_{0}),\alpha,\delta)$ represents the three-dimensional (3D) position in the sky. $\bm{x}_{\rm gw}=(d_{{\rm L},i},\alpha_{i},\delta_{i})$ represents the 3D position of the GW source. Considering the bias between the true value and the central value of the posterior distribution from the actual observation, we use the Fisher matrix to calculate the Gaussian distribution of $x_{\rm gw}$, and randomly select the value of $x_{\rm gw}$ according to this Gaussian distribution. We establish the boundary of the GW source’s localization volume based on the 3D measurement errors of $x_{\rm gw}$. The galaxies whose positions satisfying $\chi^{2}\leq 11.34$ (corresponding to $99\%$ confidence) can be considered as the potential host galaxies of the GW source. The number of potential host galaxies for a GW source is defined as $N_{\rm in}$, which is depend on the prior values of $H_{0}$. We display $N_{\rm in}$ as a function of SNR in Fig. 4, fixing $H_{0}=67.66~{}\rm km\ s^{-1}~{}\rm Mpc^{-1}$. Due to the lack of EM signals in dark sirens, $p\left(\mathcal{D}_{\mathrm{EM},i}\mid z,\alpha,\delta\right)$ can be set to a constant value Chen et al. (2018).
For the potential candidate host galaxies, $p_{0}(z,\alpha,\delta\mid H_{0})$ in Eq. (27) can be expressed as
$p_{0}\left(z,\alpha,\delta\mid H_{0}\right)=$ $f_{\rm com}p_{\text{cat}}(z,\alpha,\delta)+(1-f_{\rm com})$ (29)
×$\times p_{\text{miss}}\left(z,\alpha,\delta\mid H_{0}\right),$
@@
$f_{\rm com}=\frac{N_{\text{cat}}}{\bar{n}_{\mathrm{gal}}V_{\rm max}},$ (30)
@@
$p_{\text{cat}}(z,\alpha,\delta)=\frac{1}{N_{\text{in}}}\sum_{j=1}^{N_{\text{in}}}$ $\mathcal{N}\left(z_{j},\sigma_{z,j}\right)\delta\left(\alpha-\alpha_{j}\right)$ (31)
×$\times\delta\left(\delta-\delta_{j}\right),$
where $\mathcal{N}$ represents the Gaussian distribution with mean $z_{j}$ and the standard deviation $\sigma_{z,j}=(1+z){\sqrt{\langle{v_{\rm p}}^{2}\rangle}}/{c}$ Hogg (1999); Muttoni et al. (2023). ${\sqrt{\langle{v_{\rm p}}^{2}\rangle}}$ is the standard deviation of the radial peculiar velocity of a galaxy, set to $500~{}\rm km~{}s^{-1}$ Henriques et al. (2012). And $p_{\rm miss}$ is given by
$p_{\text{miss}}(z,\alpha,\delta\mid H_{0})\propto(1-p_{\text{com}}(z,\alpha,\delta))\frac{\mathrm{d}V_{\rm c}}{\mathrm{d}z\mathrm{d}\alpha\mathrm{d}\delta}.$ (32)
The normalization term $\gamma(H_{0})$ in Eq. (26) can be written as
$\gamma\left(H_{0}\right)=\int$ $p_{\operatorname{det}}^{\mathrm{GW}}\left({d}_{\mathrm{L}}\left(z,H_{0}\right),\alpha,\delta\right)p_{\operatorname{det}}^{\mathrm{EM}}(z,\alpha,\delta)$ (33)
×$\times p_{0}(z,\alpha,\delta)\mathrm{d}z\mathrm{d}\alpha\mathrm{d}\delta,$ (34)
where $p_{\mathrm{det}}^{\mathrm{GW}}$ and $p_{\mathrm{det}}^{\mathrm{EM}}$ represent the GW selection effect and the EM selection effect, respectively. The GW selection effect can be expressed as
$&p_{\operatorname{det}}^{\mathrm{GW}}\left(d_{\mathrm{L}}\left(z,H_{0}\right),\alpha,\delta\right)=\\ &\int_{\mathcal{D}_{\mathrm{GW}}>\mathcal{D}_{\mathrm{GW}}^{\text{th}}}p\left(\mathcal{D}_{\mathrm{GW}}\mid{d}_{\mathrm{L}}\left(z,H_{0}\right),\alpha,\delta\right)\mathrm{d}\mathcal{D}_{\mathrm{GW}},\end{aligned}$ (35)
where $\mathcal{D}^{\rm th}_{\rm GW}$ denotes the SNR threshold for GW detections and is set to 10. The EM selection effect can be expressed as $p_{\mathrm{det}}^{\mathrm{EM}}(z,\alpha,\delta)\propto\mathcal{H}\left(z_{\max}-z\right)$, where $\mathcal{H}$ is the Heaviside step function and $z_{\max}=0.1$.
@@ III Results and discussions
      $\sigma_{t}=20\ [\text{ns}],N_{\rm p}=20$       5.0       49       2.88%       174       2.44%
      10.0       63       2.42%       237       2.07%
      15.0       67       2.20%       274       1.88%
      20.0       80       2.07%       295       1.71%
      $\sigma_{t}=20\ [\text{ns}],N_{\rm p}=100$       5.0       105       1.90%       377       1.33%
      10.0       130       1.60%       421       1.05%
      15.0       138       1.45%       451       0.90%
      20.0       150       1.35%       461       0.85%
      $\sigma_{t}=20\ [\text{ns}],N_{\rm p}=200$       5.0       133       1.60%       439       1.07%
      10.0       152       1.38%       457       0.86%
      15.0       164       1.30%       481       0.73%
      20.0       173       1.21%       487       0.71%
@@
In this section, we compare the two SMBHB population models, discuss the GW parameter estimations, and present the constraint results for $H_{0}$ from various PTA detection cases. As shown in Fig. 2, the $h_{0}$, $z$, $M_{\rm BHB}$, and $f_{0}$ distributions of SMBHBs simulated from the two population models are essentially similar. Nevertheless, the numbers of simulated SMBHBs for the two population models are 490 and 213, respectively, showing that the C21 model has a noticeably larger number of SMBHBs than the M17 model. We consider the 10-year observations based on the C21 model as the representative example of PTA observations for the following discussions.
In Fig. 5, we present the distributions of $\rho$, $\Delta\Omega$, and $\Delta d_{\rm L}/d_{\rm L}$ based on the 20, 100, and 200 MSPs with $\sigma_{t}=20$ ns. Note that the calculations are based on 213 (M17 model) and 490 (C21 model) simulated SMBHBs at $z<0.1$. As shown in the left panel of Fig. 5, fewer available pulsars result in lower SNRs, thus showing worse localization abilities. We note that the localization capability with 100 MSPs is much better than that of 20 MSPs, but 200 MSPs are slightly better than 100 MSPs. This indicates that 100 MSPs seem to be sufficient and more economical, as further increasing the number of pulsars does not significantly enhance the localization ability.
@@
Figure 6 illustrates the relative errors for $H_{0}$ as a function of $T_{\rm span}$. Table 1 shows the relative error of $H_{0}$, defined as $\varepsilon(H_{0})=\sigma(H_{0})/H_{0}$. We observe that the PTAs with $N_{\rm p}=20$ give the poorest measurement precision for $H_{0}$. Based on the C21 model, the constraint precision of $H_{0}$ is 2.44% for a 5-year observation, improving to 1.71% over a 20-year observation. These values are worse than the constraint precision obtained from current distance ladder measurements. Increasing the number of MSPs to 100 significantly enhances the precision of $H_{0}$ compared to using 20 MSPs. With a 5-year observation, the constraint precision of $H_{0}$ is 1.33% and can be improved to 0.85% with a 20-year observation; both values are close to or better than 1%. Furthermore, monitoring 200 MSPs provides even better constraints on $H_{0}$, achieving precisions ranging from 1.07% to 0.71% over 5 to 20 years of observations.
@@
Figure 7 presents the constraints on $H_{0}$ obtained from individual GW events as well as the combined results. The colored lines represent the constraints from individual events and the black line represents the combined result of all events. There is a notable discrepancy between the two SMBHB population models in terms of the $H_{0}$ posterior distributions given by the individual events. In the M17 model, there are several events whose SNRs far exceed the others, but in the C21 model, there is no significant difference in the SNRs given by all the events. This discrepancy can also be seen in Fig. 2: Several SMBHBs (represented by the orange circles) in the M17 model have significantly higher masses and GW strains than the others. This is due to the fact that the mechanisms of the two models in predicting the SMBHB populations are substantially different. The M17 model predicts the masses of SMBHBs based on the apparent magnitude of the galaxies in the galaxy catalog. A few of these galaxies have significantly larger masses, and the SMBHBs inhabiting them could produce the GW signals with notably higher SNRs than the other SMBHBs. However, for the C21 model, its SMBHB distribution is derived from the SGWB signal, which is the overall superposition of many GW events and does not reflect the information of individual sources. As a result, the SMBHB distribution predicted by the C21 model is statistical rather than individualistic. Therefore, in the C21 model, no event exhibits a SNR that is significantly higher than other events.
@@
Compared with the dark siren analysis in Ref. Wang et al. (2022d), under the same detection conditions and population models, the constraint precision of $H_{0}$ is improved in this paper. This is mainly due to the fact that we simulate SMBHBs over a broader range of redshift ($0.10<z<$ 0.1), resulting in a larger number of GW sources available as dark sirens. We further consider the effect of incompleteness for galaxies at $z>0.04$. The authors of Ref. Wang et al. (2022d) also indicated some advantages of nHz GW standard sirens over the standard sirens in other frequency bands. Here, we want to emphasize another advantage: Since SMBHBs are all hosted in supermassive galaxies, we can directly exclude those galaxies with low masses when using the dark siren method, and only need to consider the supermassive galaxies within the localization volume of the GW source. Compared with the stellar-mass dark sirens, it can greatly reduces the computational costs and the uncertainties caused by spurious host galaxies in Bayesian inference.
IV Conclusion

### v2

@@ 2.3 GW detection and Fisher analysis
$\Delta{d_{\mathrm{L}}}$ $=\sqrt{\left(\Delta{d_{\mathrm{L}}^{\rm inst}}\right)^{2}+\left(\Delta{d_{\mathrm{L}}^{\rm lens}}\right)^{2}}.$ (2.24)
The distances of MSPs are expected to be measured via the timing parallax method in the SKA era. We estimate the uncertainties of MSPs’ distances by using Eq. (18) in Lee et al. [142],
$\sigma_{d_{\rm p}}\simeq\frac{2.34}{\cos^{2}\beta_{\rm p}}\left(\frac{N}{100}\right)^{-1/2}\left(\frac{d_{\rm p}}{1~{}\mathrm{kpc}}\right)^{2}\left(\frac{\sigma_{t}}{10~{}\mathrm{ns}}\right)~{}\mathrm{pc},$ (2.25)
where $\beta_{\rm p}$ represents the ecliptic latitudes of MSPs. In the Fisher matrix, we take into account the priors of MSPs’ distances using Eq. (2.20). In this context, $P(d_{\rm p})$ is a Gaussian distribution with mean $d_{\rm p}$ and the standard deviation $\sigma_{d_{\rm p}}$. Under this assumption, the derivatives of GW frequencies can be obtained, and the chirp masses of the SMBHBs can be further measured. This enables us to break the degeneracy between ${M_{\rm c}}$ and $d_{\rm L}$, and therefore precisely infer $d_{\rm L}$ from GW amplitudes. The distributions of the $d_{\rm L}$ relative errors and the angular resolutions of simulated SMBHBs versus SNR based on 20, 100, and 200 MSPs with $\sigma_{t}$ = 20 $\rm ns$ and 10-year observation are displayed in figure 3. The sky locations of MSPs are randomly selected from the Australia Telescope National Facility pulsar catalog [182].
@@
2.4 Inference of the Hubble constant
@@
$H(z)=H_{0}\sqrt{\Omega_{\rm m}(1+z)^{3}+1-\Omega_{\rm m}},$ (2.26)
@@
$d_{\rm L}(z)=c(1+z)\int_{0}^{z}\frac{1}{H(z^{\prime})}\mathrm{d}z^{\prime},$ (2.27)
where $c$ is the speed of light in vacuum. Throughout this work, we adopt the flat $\Lambda$CDM model to generate the dark siren data with the best-fit values of $\Omega_{\rm m}=0.3111$ and $H_{0}=67.66\rm~{}km~{}s^{-1}~{}Mpc^{-1}$ given by the Planck 2018 TT, TE, EE+lowE+lensing+BAO data [3]. Based on the flat $\Lambda$CDM model, we utilize the GW signals from SMBHBs as dark sirens to infer $H_{0}$, fixing the other cosmological parameters to the Planck 2018 results.
@@
$p\left(H_{0}\mid\mathcal{D}_{\mathrm{GW}},\mathcal{D}_{\mathrm{EM}}\right)=\frac{\prod_{i=1}^{N_{\mathrm{GW}}}p\left(\mathcal{D}_{\mathrm{GW},i},\mathcal{D}_{\mathrm{EM},i}\mid{H_{0}}\right)p({H_{0}})}{\gamma(H_{0})},$ (2.28)
@@
$p\left(\mathcal{D}_{\mathrm{GW},i},\mathcal{D}_{\mathrm{EM},i}\mid H_{0}\right)$ (2.29)
@@
$=\int p\left(\mathcal{D}_{\mathrm{GW},i}\mid{d}_{\rm{L}},\alpha,\delta\right)p\left(\mathcal{D}_{\mathrm{EM},i}\mid z,\alpha,\delta\right)p\left({d}_{\rm{L}}\mid z,H_{0}\right)p_{0}\left(z,\alpha,\delta\mid H_{0}\right)\mathrm{d}{d}_{\rm{L}}\mathrm{d}\alpha\mathrm{d}\delta\mathrm{d}z$
$=\int p\left(\mathcal{D}_{\mathrm{GW},i}\mid{d}_{\rm{L}},\alpha,\delta\right)p\left(\mathcal{D}_{\mathrm{EM},i}\mid z,\alpha,\delta\right)\delta\left.(d_{\rm{L}}-\hat{d}_{\rm{L}}\left(z,H_{0}\right)\right.)p_{0}\left(z,\alpha,\delta\mid H_{0}\right)\mathrm{d}d_{\rm{L}}\mathrm{d}\alpha\mathrm{d}\delta\mathrm{d}z$
$=\int p\left.(\mathcal{D}_{\mathrm{GW},i}\mid\hat{d}_{\rm L}\left(z,H_{0}\right),\alpha,\delta\right.)p\left(\mathcal{D}_{\mathrm{EM},i}\mid z,\alpha,\delta\right)p_{0}\left(z,\alpha,\delta\mid H_{0}\right)\mathrm{d}\alpha\mathrm{d}\delta\mathrm{d}z.$
@@
$p\left.(\mathcal{D}_{\mathrm{GW},i}\mid\hat{d}_{\mathrm{L}}\left(z,H_{0}\right),\alpha,\delta\right.)\propto e^{-\chi^{2}/2},$ (2.30)
with $\chi^{2}=(x-x_{\rm gw})^{T}\mathcal{C}^{-1}(x-x_{\rm gw})$, where $\mathcal{C}$ is the $3\times 3$ covariance matrix only relevant to $(d_{\rm{L}},\alpha,\delta)$, obtained from the Fisher information matrix. $x=(\hat{d}_{\rm{L}}(z,H_{0}),\alpha,\delta)$ represents the three-dimensional (3D) position in the sky. $\bm{x}_{\rm gw}=(d_{{\rm L},i},\alpha_{i},\delta_{i})$ represents the 3D position of the GW source. Considering the bias between the true value and the posterior median from the actual observation, we use the Fisher matrix to determine the Gaussian distribution of $x_{\rm gw}$ and randomly sample $x_{\rm gw}$ from this distribution. We establish the boundary of the GW source’s localization volume based on the 3D measurement errors of $x_{\rm gw}$. The galaxies whose positions satisfying $\chi^{2}\leq 11.34$ (corresponding to $99\%$ confidence) can be considered as the potential host galaxies of the GW source. The number of potential host galaxies for a GW source is defined as $N_{\rm in}$, which is dependent on the prior values of $H_{0}$. We display $N_{\rm in}$ as a function of SNR in figure 4, fixing $H_{0}=67.66~{}\rm km\ s^{-1}~{}\rm Mpc^{-1}$. Due to the lack of EM signals in dark sirens, $p\left(\mathcal{D}_{\mathrm{EM},i}\mid z,\alpha,\delta\right)$ can be set to a constant value [89].
For the potential candidate host galaxies, $p_{0}(z,\alpha,\delta\mid H_{0})$ in Eq. (2.29) can be expressed as
$p_{0}\left(z,\alpha,\delta\mid H_{0}\right)=$ $f_{\rm com}p_{\text{cat}}(z,\alpha,\delta)+(1-f_{\rm com})p_{\text{miss}}\left(z,\alpha,\delta\mid H_{0}\right),$ (2.31)
@@
$f_{\rm com}=\frac{N_{\text{cat}}}{\bar{n}_{\mathrm{gal}}V_{\rm max}},$ (2.32)
@@
$p_{\text{cat}}(z,\alpha,\delta)=\frac{1}{N_{\text{in}}}\sum_{j=1}^{N_{\text{in}}}$ $\mathcal{N}\left(z_{j},\sigma_{z,j}\right)\delta\left(\alpha-\alpha_{j}\right)\delta\left(\delta-\delta_{j}\right),$ (2.33)
where $\mathcal{N}$ represents the Gaussian distribution with mean $z_{j}$ and the standard deviation $\sigma_{z,j}=(1+z){\sqrt{\langle{v_{\rm p}}^{2}\rangle}}/{c}$ [183, 103]. ${\sqrt{\langle{v_{\rm p}}^{2}\rangle}}$ is the standard deviation of the radial peculiar velocity of a galaxy, set to $500~{}\rm km~{}s^{-1}$ [184]. The expression for $p_{\rm miss}$ is given by
$p_{\text{miss}}(z,\alpha,\delta\mid H_{0})\propto(1-p_{\text{com}}(z,\alpha,\delta))\frac{\mathrm{d}V_{\rm c}}{\mathrm{d}z\mathrm{d}\alpha\mathrm{d}\delta}.$ (2.34)
The normalization term $\gamma(H_{0})$ in Eq. (2.28) can be written as
$\gamma\left(H_{0}\right)=\int$ $p_{\operatorname{det}}^{\mathrm{GW}}\left({d}_{\mathrm{L}}\left(z,H_{0}\right),\alpha,\delta\right)p_{\operatorname{det}}^{\mathrm{EM}}(z,\alpha,\delta)p_{0}(z,\alpha,\delta)\mathrm{d}z\mathrm{d}\alpha\mathrm{d}\delta,$ (2.35)
where $p_{\mathrm{det}}^{\mathrm{GW}}$ and $p_{\mathrm{det}}^{\mathrm{EM}}$ represent the GW detection probability and the EM detection probability, respectively. The GW detection probability can be expressed as
$&p_{\operatorname{det}}^{\mathrm{GW}}\left(d_{\mathrm{L}}\left(z,H_{0}\right),\alpha,\delta\right)=\int_{\mathcal{D}_{\mathrm{GW}}>\mathcal{D}_{\mathrm{GW}}^{\text{th}}}p\left(\mathcal{D}_{\mathrm{GW}}\mid{d}_{\mathrm{L}}\left(z,H_{0}\right),\alpha,\delta\right)\mathrm{d}\mathcal{D}_{\mathrm{GW}},\end{aligned}$ (2.36)
where $\mathcal{D}^{\rm th}_{\rm GW}$ denotes the SNR threshold for GW detections and is set to 10. The EM detection probability can be expressed as $p_{\mathrm{det}}^{\mathrm{EM}}(z,\alpha,\delta)\propto\mathcal{H}\left(z_{\max}-z\right)$, where $\mathcal{H}$ is the Heaviside step function and $z_{\max}=0.1$.
@@
$\sigma_{t}=20\ [\text{ns}],N_{\rm p}=20$ 5.0 49 7.75% 174 3.56%
10.0 63 4.49% 237 2.91%
15.0 67 3.10% 274 2.27%
20.0 80 2.58% 295 1.82%
$\sigma_{t}=20\ [\text{ns}],N_{\rm p}=100$ 5.0 105 2.21% 377 1.48%
10.0 130 1.78% 421 1.18%
15.0 138 1.54% 451 0.97%
20.0 150 1.44% 461 0.91%
$\sigma_{t}=20\ [\text{ns}],N_{\rm p}=200$ 5.0 133 1.72% 439 1.11%
10.0 152 1.36% 457 0.88%
15.0 164 1.18% 481 0.77%
20.0 173 1.11% 487 0.68%
@@
3 Results and discussions
In this section, we compare the two SMBHB population models, discuss the GW parameter estimations, and present the constraint results for $H_{0}$ from various PTA detection cases. As shown in figure 2, the $h_{0}$, $z$, $M_{\rm BHB}$, and $f_{0}$ distributions of SMBHBs simulated from the two population models are essentially similar. Nevertheless, the numbers of simulated SMBHBs for the two population models are 490 and 213, respectively, showing that the C21 model has a noticeably larger number of SMBHBs than the M17 model. We consider the 10-year observations based on the C21 model as the representative example of PTA observations for the following discussions.
In figure 5, we present the distributions of $\rho$, $\Delta\Omega$, and $\Delta d_{\rm L}/d_{\rm L}$ based on the 20, 100, and 200 MSPs with $\sigma_{t}=20$ ns. Note that the calculations are based on 213 (M17 model) and 490 (C21 model) simulated SMBHBs at $z<0.1$. As shown in the left panel of figure 5, fewer available MSPs result in lower SNRs, thus showing worse localization abilities. We note that the localization capability with 100 MSPs is much better than that of 20 MSPs, but 200 MSPs are slightly better than 100 MSPs. This indicates that 100 MSPs seem to be sufficient and more economical, as further increasing the number of MSPs does not significantly enhance the localization ability.
@@ 3 Results and discussions
Figure 6 illustrates the relative errors for $H_{0}$ as a function of $T_{\rm span}$. Table 1 shows the relative error of $H_{0}$, defined as $\varepsilon(H_{0})=\sigma(H_{0})/H_{0}$. We observe that the PTAs with $N_{\rm p}=20$ give the poorest measurement precision for $H_{0}$. Based on the C21 model, the constraint precision of $H_{0}$ is 3.56% for a 5-year observation, improving to 1.82% over a 20-year observation. These values are worse than the constraint precision obtained from current distance ladder measurements. Increasing the number of MSPs to 100 significantly enhances the precision of $H_{0}$ compared to using 20 MSPs. With a 5-year observation, the constraint precision of $H_{0}$ is 1.48% and can be improved to 0.91% with a 20-year observation; both values are close to or better than 1%. Furthermore, monitoring 200 MSPs provides even better constraints on $H_{0}$, achieving precisions ranging from 1.11% to 0.68% over 5 to 20 years of observations. Considering that SKA is expected to begin operations around 2030, our optimistic estimate ($\sigma_{t}=20$ ns, $N_{\rm p}=200$) suggests that PTA dark sirens can achieve a measurement precision of approximately 1% for $H_{0}$ over a 10-year observation, which would be around 2040. Nevertheless, under a more conservative noise assumption and considering the impact of the SGWB, the first results are more likely to be obtained after 2050. In figure 6, we plot the current $H_{0}$ values obtained from the distance ladder measurements and the Planck CMB data. It should be noted that when the precision of the PTA dark siren method in measuring $H_{0}$ reaches its expected level in the future, the measurement accuracy of the distance ladder method will be also significantly enhanced. Although it is difficult for the PTA dark siren method to surpass the distance ladder method in terms of precisely measuring $H_{0}$, it can still provide assistance in resolving the Hubble tension as an independent late-universe probe.
It is worth noting that our results are influenced by the value of $h_{c}$. Since $h_{c}^{2}\propto\Phi_{\rm BHB,0}$, a larger $h_{c}$ corresponds to a greater inferred number of SMBHBs. This increases the number of detectable SMBHBs serving as dark sirens, and further enhances the measurement precision of $H_{0}$. Compared with the NANOGrav 12.5-yr data, the $h_{c}$ value inferred from the NANOGrav 15-yr dataset suggests a larger SMBHB population. However, the $h_{c}$ values reported by NANOGrav [130], EPTA+InPTA [132], PPTA [134] and CPTA [136] in 2023 are generally consistent. Therefore, using different $h_{c}$ values provided by various PTA projects will not alter our main conclusions.
@@
Figure 7 presents the constraints on $H_{0}$ obtained from individual GW events as well as the combined results. The colored lines represent the constraints from individual events and the black line represents the combined result of all events. There is a notable discrepancy between the two SMBHB population models in terms of the $H_{0}$ posterior distributions given by the individual events. In the M17 model, there are several events whose SNRs far exceed the others, but in the C21 model, there is no significant difference in the SNRs given by all the events. This discrepancy can also be seen in figure 2: Several SMBHBs (represented by the orange circles) in the M17 model have significantly higher masses and GW strains than the others. This is due to the fact that the mechanisms of the two models in predicting the SMBHB populations are substantially different. The M17 model predicts the masses of SMBHBs based on the apparent magnitude of the galaxies in the galaxy catalog. A few of these galaxies have significantly larger masses, and the SMBHBs inhabiting them could produce the GW signals with notably higher SNRs than the other SMBHBs. However, for the C21 model, its SMBHB distribution is derived from the SGWB signal, which is the overall superposition of many GW events and does not reflect the information of individual sources. As a result, the SMBHB distribution predicted by the C21 model is statistical rather than individualistic. Therefore, in the C21 model, no event exhibits a SNR that is significantly higher than other events.
PTA dark sirens primarily rely on low-redshift ($z<0.1$) GW sources and galaxy catalogs. Within this range, the distance-redshift relation can be approximated by Hubble’s law and has limited dependence on late-time cosmological extended models with new physics, such as time-varying dark energy [185, 186], interacting dark energy [187, 26], phantom dark energy [188, 189], decaying dark matter [190, 191], and spatial curvature [192, 193]. Consequently, these additional parameters cannot be effectively constrained by the low-redshift PTA dark sirens. If these extended models are considered, the additional introduced parameters would exhibit degeneracy with $H_{0}$, thereby slightly diminishing the measurement precision of $H_{0}$.
@@
Compared with the dark siren analysis in Wang et al. [149], under the same detection conditions and population models, the constraint precision of $H_{0}$ is improved in this paper. This is mainly due to the fact that we simulate SMBHBs over a broader range of redshift ($0.10<z<$ 0.1), resulting in a larger number of GW sources available as dark sirens. We further consider the effect of incompleteness for galaxies at $z>0.04$. Wang et al. [149] also indicated some advantages of nHz GW standard sirens over the standard sirens in other frequency bands. Here, we want to emphasize another advantage: Since SMBHBs are all hosted in supermassive galaxies, we can directly exclude those galaxies with low masses when using the dark siren method, and only need to consider the supermassive galaxies within the localization volume of the GW source. Compared with the stellar-mass dark sirens, it can greatly reduce the computational costs and the uncertainties caused by spurious host galaxies in Bayesian inference.
4 Conclusion